OpenStudy (anonymous):

14/4m = 16/5m+9 (It's fractions not division :3) 14/4m = 16/5m+9 (It's fractions not division :3) @Mathematics

OpenStudy (anonymous):

7/2m=16/5m+9 = 7/2m-16/5m =9 =(35-32)/10m=9 3 = 90m m=1/30